# Phase behavior of Cacio and Pepe sauce

### v2

@@ Abstract
“Pasta alla Cacio e pepe” is a traditional Italian dish made with pasta, pecorino cheese, and pepper. Despite its simple ingredient list, achieving the perfect texture and creaminess of the sauce can be challenging. In this study, we systematically explore the phase behavior of Cacio and pepe sauce, focusing on its stability at increasing temperatures for various proportions of cheese, water, and starch. We identify starch concentration as the key factor influencing sauce stability, with direct implications for practical cooking. Specifically, we delineate a regime where starch concentrations below 1% (relative to cheese mass) lead to the formation of system-wide clumps, a condition determining what we term the “Mozzarella Phase” and corresponding to an unpleasant and separated sauce. Additionally, we examine the impact of cheese concentration relative to water at a fixed starch level, observing a lower critical solution temperature that we theoretically rationalized by means of a minimal effective free-energy model. Finally, we present a scientifically optimized recipe based on our findings, enabling a consistently flawless execution of this classic dish.
@@ I Introduction
On several occasions, pasta has been a source of inspiration for physicists [1]. The observation that spaghetti always break up into three or more fragments, but never in two halves, puzzled even Richard Feynman himself, and the explanation of this intriguing phenomenon earned Audoly and Neukirch the Ig Nobel Prize [2]. Pasta packaging offers a natural framework to study separation upon shaking [3] and inspired the design of “morphing flat pasta” [4]. Furthermore, the deformation and swelling behavior of various pasta varieties upon cooking has been experimentally and theoretically investigated [5, 6, 7, 8]. In particular, Hwang et al. found that the stick length of spaghetti correlates with how much the pasta is cooked, proposing an experiment to assess if pasta is “al dente”  [5]. Lastly, the study of polymer rings and topological glasses has inspired a pasta prototype, a kind of closed-loop linguine that the authors have named “Anelloni” [9].
In addition to inspiring studies on its mechanical properties and structural behavior, pasta water (a mixture of starch and water) also exhibits interesting physical properties. These properties, which are central to both cooking and scientific exploration, include intriguing rheological phenomena and transitions such as non-Newtonian behavior and gelation [10]. Adding 1 part of water to 1.5–2 parts of corn starch is the most common way to get a non-newtonian fluid, sometimes called oobleck [11]. Furthermore, upon heating, the starch and water solution undergoes a gelation transition [12, 13]. Starch-enriched water can be also fundamental in stabilizing emulsions, such as in the classic “spaghetti aglio e olio” (pasta dish simply made with garlic and oil), where it helps to form a creamy sauce by preventing the formation of suspended oil droplets.
In the culinary realm, there are many other examples where the thermodynamics of mixtures plays a crucial role  [14]. Phase separation often controls food texture [15], for example in emulsions like mayonnaise, salad dressing, and other sauces [16]. In such emulsions, the homogeneous state prepared by blending or shaking the mixture is metastable and evolves towards a thermodynamic stable state composed of oily droplets, with undesired consistency. Thus, emulsions must be stabilized to remain homogeneous longer, i.e. ripening must be slowed down [17]. Delaying ripening is crucial in many other examples including chocolate [18] and the ice cream industry  [19], where phase separation causes product degradation.
Moving to beverages, a famous example is the ouzo-effect, occurring during the dilution of anise-based alcohols, such as ouzo and pastis [20, 21]. When liquid water or ice is poured into these liquors, the solution becomes opaque due to the formation of oil droplets with sizes comparable to the wavelength of light. These droplets remain stable likely due to a trapping mechanism [22]. Similar droplets were recently found in limoncello, a famous Italian liquor [23].
More recently, phase separation physics has received a new wave of enthusiasm in biochemistry, caused by the discovery of phase-separated compartments in the cell cytoplasm [24, 25] and mixtures of oligomers mimicking the primordial soup [26].
Building upon the recent theoretical and experimental advances in phase separation, in this paper, we investigate the phase behavior of Cacio and pepe pasta sauce. Cacio and pepe (literally “cheese and pepper”) is a traditional recipe from Lazio, a region in central Italy. It consists of tonnarelli noodles served in a cream of pecorino cheese, pepper, and starch-enriched water, see Fig. 1 a). Its origin is commonly associated with the long travels of shepherds, who had to stuff their saddlebags with hypercaloric ingredients. Pecorino cheese was ideal due to its extraordinary shelf life, black pepper was used to stimulate heat receptors, and homemade spaghetti provided the carbohydrate intake.
@@
Figure 1: Cacio and pepe pasta sauce consists of pecorino cheese, pepper and starch-enriched water a) Tonnarelli noodles garnished with a mixture of pecorino cheese and starch-enriched water, and a sparkle of pepper. b) Snapshots of the mixture that constitutes the base for the pasta sauce, i.e. cheese and water with different amounts of starch, at different temperatures. In particular, we compare the effect of: water alone; pasta water that retains some starch (obtained by cooking $100$ g of pasta in $1$ liter of water); and pasta water “risottata”, i.e. pasta water heated in a pan to let the water evaporate (until reducing its total weight by three times) and starch gets concentrated. As the starch concentration increases, cheese clusters decrease in size and occur at higher temperatures. The region here named “Mozzarella phase” is characterized by huge mozzarella-like clumps of cheese suspended in water, resulting from extreme protein aggregation on heating.
@@
Despite the short list of ingredients, preparing this dish requires extra care. The most delicate step is the mixing of starch-enriched water with grated cheese. One starts cooking pasta noodles in boiling water, letting them release starch, then extracts part of the water and starch solution. An essential procedure is to wait some time before mixing water and cheese, to let the water cool down. This is because, at high temperatures, cheese proteins can either form clumps upon denaturation or simply aggregate [27, 28], therefore ruining the sauce, see Fig. 1 b). But temperature is not the only physical parameter that must be carefully controlled: protein aggregation and denaturation are concentration-dependent processes, thus mixing the right amount of cheese, water, and starch is essential to avoid protein aggregates. In the absence of starch, for example, cheese in hot water forms huge clumps at a temperature around $65\degree\text{C}$, see the first row of Fig. 1 b). If cheese is mixed with pasta water, which contains small amounts of starch, the emergence of clumps is reduced, and large protein aggregates are found at higher temperatures, see Fig. 1 b), second row. Pasta water can be “risottata”, i.e. collected and heated in a pan, so that some water evaporates and the starch is concentrated. If cheese is mixed with pasta water “risottata”, the presence of clumps is almost negligible, see the last row in Fig. 1 b).
@@
To overcome clump formation and achieve the perfect Cacio and pepe, in this work, we characterize the phase behavior of the solution containing water, starch, and cheese, by systematically investigating the role of each one of these components. We achieve this using common kitchen tools, ensuring that our results are easily reproducible not only by scientists in the lab but also by culinary enthusiasts. The paper is organized as follows: in section II, we introduce the experimental setup and discuss the role of starch in the mixture phase phenomenology. In section III, we fix the starch percentage and vary the cheese amount. Section IV discusses a minimal theoretical model that recapitulates our experimental finding. Finally, in section V we propose a recipe for the perfect Cacio and pepe.
II Increasing starch mitigates the formation of protein clumps
@@
Figure 2: Starch mitigates protein clumps. a) Phase diagram of the sauce state as a function of the starch percentage with respect to the water and temperature in Celsius. Each box contains a snapshot of the sauce mixture taken during the experiment and its contour reflects the mean elongation of the corresponding sample via the color map shown on the left. The “mozzarella phase” indicates a region of the phase diagram where the cheese in the sample forms a clump of a size comparable to the one of the system. b) The same phase diagram as in a) after Gaussian smoothening is represented as a discrete heatmap with color mapping the mean elongation of the sample. c) Kernel regression smoothening of the phase diagram in a) to obtain a continuous map. The color map on the right refers to panels b) and c).
As discussed above, one of the most problematic aspects of making Cacio and Pepe is the appearance of large aggregates of cheese (Mozzarella phase). These clumps make the sauce inhomogeneous, giving it a gummy, unpleasant consistency. Fig. 1b) shows how, in the absence of starch in the mixture, increasing the temperature leads to the abrupt appearance of a giant dense phase. From a culinary point of view, this represents a problem because small temperature variations can completely compromise the recipe’s outcome. The addition of pasta water, which contains small amounts of starch, mitigates the suddenness of the Mozzarella phase emergence (Fig. 1b), second row), and increasing starch concentration even more (as in the pasta water "risottata") suppresses the mozzarella Phase altogether (Fig. 1b), third row). Starch thus seems to stabilize the homogeneous mixture over a large range of temperatures, allowing for less stringent temperature control during preparation.
To test this hypothesis, we conducted experiments at a fixed cheese-to-water ratio while varying the starch concentration in the water and the temperature of the system We prepared the starch-enriched water, dissolving dry corn starch into water at ambient temperature, targeting the weight percentages shown on the bottom axis of Fig. 2 a). The mixture was heated on a stovetop to gelation, a transition marked by a noticeable increase in viscosity and opacity. Afterward, we allowed the starch-enriched water to cool to room temperature to prevent excessive heating of the cheese-water mixture during blending. Cheese and starch-enriched water were combined in equal weights and thoroughly blended with a mixer. The resulting mixture was placed in a controlled heat bath — a pot of water maintained at a constant temperature using a sous vide cooker device. The mixture’s temperature was further monitored with an external thermometer positioned inside it, and samples were collected upon reaching thermal equilibrium, as indicated by a stable mixture temperature. Each sample was spooned onto a Petri dish and photographed. In the Supplemental Material, we report on the experimental apparatus built to heat the mixture and collect images, along with a detailed explanation of the measurement protocol. Each mixture was imaged systematically increasing the heat bath temperature by $5$C at each sampling and the evaporated water was replenished at each temperature step. Experiments were performed at different starch percentages to generate the images forming the phase diagram in Fig. 2 a). To characterize the mixture phase behavior, we approximated the cheese clusters as detected by quantile thresholding as ellipses and quantified their elongation in each sample (see Supplemental Material for details on data analysis). As an order parameter, we considered the mean cluster elongation since the phenomenon of protein aggregation is often referred to, in Italian kitchen jargon, as cheese “filatura”, i.e., the creation of stripes of cheese. We remark that other measures characterizing clusters’ size (or relative size) gave qualitatively similar results. Samples with larger, better separated clumps resulted in higher mean elongation values, while smoother mixtures corresponded to lower elongation values. A colored frame is used in Fig. 2 a) to associate each image to the respective value of the order parameter, measured in millimeters.
From these measurements, we constructed a phase diagram by plotting mean elongation as a function of starch concentration and temperature (see Fig.2 b)). These data were Gaussian-smoothed to create a continuous phase diagram (see Fig.2 c)). The final results reveal a clear quantitative impact of starch concentration: higher starch contents shift the transition to clumpier, less homogeneous mixtures to higher temperatures. Moreover, mixtures with low starch concentrations exhibited larger aggregate sizes, while higher starch contents reduced aggregate size and delayed clump formation. This finding corroborates the culinary insight that starch in pasta water stabilizes the Cacio and pepe sauce.
Overall, these experiments provided a quantitative picture of the starch’s role in the cheese-water mixture phase behavior. It reduces phase separation, delays the onset of clump formation to higher temperatures, and minimizes the size of aggregates, making the sauce less sensitive to mistakes in temperature control during preparation.
@@
Figure 3: Protein amount tunes aggregate formation. a) Phase diagram of the sauce made by combining $50$ g of cheese with a varying amount of starch-enriched water (here expressed as a mass percentage) as a function of the mixture temperature. The starch percentage is fixed to $1$. Each box is colored with respect to the mean elongation of cheese clumps. b) Same phase diagram as in a) after Gaussian smoothening, where each point is colored according to the map of the mean clump elongation (color map on the right). c) Phase diagram expressed as a function of protein percentage, which is the component that leads to aggregation. Kernel regression smoothening has been applied to obtain a continuous diagram. The dashed gray line indicates the isoline of median elongation, while the dashed black line represents the parabolic fit separating homogeneous from clumped domains.
III Cheese concentration tunes the emergence of protein clumps
@@
Having described the sauce mixture as a function of the starch, we ought to find the mixture’s degree of phase separation by varying the respective percentages of water and cheese. To this end, we fixed the percentage of starch in the water to $1$, a value potentially enabling the appearance of rich phase behavior. Following the procedure outlined above, we performed experiments at different percentages of starch-enriched water with respect to cheese. In Fig. 3 a), we show the resulting images, implementing a colored frame to highlight the mean elongation, the parameter employed to quantify the degree of phase separation. As expected, by fixing the concentration of starch to 1 %, we avoid the “Mozzarella phase” emerging in some of the previous experiments (see Fig. 2). From the phase diagram, we note a smooth binodal region that separates a well-mixed to a phase-separated phase, described by the mean elongation of the clumps. In Fig. 3 b), we construct a phase diagram as a function of water percentage and temperature. By obtaining the effective amount of proteins from the water percentage, in Fig. 3 c), we apply a Gaussian smoothing to obtain a continuous phase diagram as a function of protein mass fraction and temperature. This approach enables the identification of the region where clumps appear as a function of the control parameters. The binodal line separating the phase-separated domain from the homogenous one has been identified with the isoline of mean elongation corresponding to the midpoint between the maximum and minimum observed values, after smoothing. The shape of the binodal is a paraboloid that can be fitted with a simple quadratic functional form. The minimum of the parabola lies slightly below the protein mass fraction value $0.134$, achieved with water and cheese in 1:1 proportion. A parabola with positive curvature signals that both lower and higher values of protein mass fraction correspond to well-mixed sauces even at high temperatures.

### v3

# Phase behavior of Cacio e Pepe sauce

@@ Abstract
“Pasta alla Cacio e pepe” is a traditional Italian dish made with pasta, pecorino cheese, and pepper. Despite its simple ingredient list, achieving the perfect texture and creaminess of the sauce can be challenging. In this study, we systematically explore the phase behavior of Cacio e pepe sauce, focusing on its stability at increasing temperatures for various proportions of cheese, water, and starch. We identify starch concentration as the key factor influencing sauce stability, with direct implications for practical cooking. Specifically, we delineate a regime where starch concentrations below 1% (relative to cheese mass) lead to the formation of system-wide clumps, a condition determining what we term the “Mozzarella Phase” and corresponding to an unpleasant and separated sauce. Additionally, we examine the impact of cheese concentration relative to water at a fixed starch level, observing a lower critical solution temperature that we theoretically rationalized by means of a minimal effective free-energy model. We further analyze the effect of a less traditional stabilizer, trisodium citrate, and observe a sharp transition from the Mozzarella Phase to a completely smooth and stable sauce, in contrast to starch-stabilized mixtures, where the transition is more gradual. Finally, we present a scientifically optimized recipe based on our findings, enabling a consistently flawless execution of this classic dish.
@@ I Introduction
Figure 1: Cacio e pepe pasta sauce consists of pecorino cheese, pepper and starch-enriched water (a) Noodles with an emulsion of pecorino cheese and starch-enriched water, garnished with freshly ground black pepper. (b) Snapshots of the mixture that constitutes the base for the pasta sauce, i.e. cheese and water with different amounts of starch, at different temperatures. In particular, we compare the effect of: water alone; pasta water that retains some starch (obtained by cooking $100$ g of pasta in $1$ liter of water); and pasta water “risottata”, i.e. pasta water heated in a pan to let the water evaporate (until reducing its total weight by three times) and starch gets concentrated. As the starch concentration increases, cheese aggregates decrease in size and occur at higher temperatures. The region here named “Mozzarella Phase” is characterized by huge mozzarella-like clumps of cheese suspended in water, resulting from extreme protein aggregation on heating.
On several occasions, pasta has been a source of inspiration for physicists [1]. The observation that spaghetti always break up into three or more fragments, but never in two halves, puzzled even Richard Feynman himself, and the explanation of this intriguing phenomenon earned Audoly and Neukirch the Ig Nobel Prize [2]. Pasta packaging offers a natural framework to study separation upon shaking [3] and inspired the design of “morphing flat pasta” [4]. Furthermore, the deformation and swelling behavior of various pasta varieties upon cooking has been experimentally and theoretically investigated [5, 6, 7]. Lastly, analogies with pasta shapes proved useful in different physics fields, from polymer rings to neutron stars  [8, 9].
Pasta water also exhibits fascinating physical properties, primarily due to its starch content. When heated, starch-water solutions undergo a gelatinization transition [10, 11], altering their viscosity and structural properties. Additionally, mixing water with 1.5–2 parts of corn starch is a well-known method for creating a non-Newtonian fluid, often referred to as oobleck [12]. Starch-enriched water also plays a crucial role in stabilizing emulsions, as seen in the classic dish “spaghetti aglio e olio”, where it helps form a smooth, creamy sauce by preventing the separation of oil into suspended droplets.
In the culinary realm, many other examples highlight the importance of emulsions containing droplets formed via phase separation [13]. Phase separation often plays a crucial role in determining food texture [14], as seen in mayonnaise, salad dressings, and various sauces [15]. In such emulsions, the homogeneous state achieved by blending or shaking is metastable and naturally evolves toward a thermodynamically stable state characterized by the formation of oily droplets, often resulting in an undesirable consistency. To maintain a uniform texture, emulsions must be stabilized to slow down this ripening process [16]. The need to delay phase separation extends beyond emulsions to other food products, such as chocolate [17] and ice cream [18], where unwanted phase separation leads to textural degradation and reduced shelf life. Droplet physics plays an important role in beverages as well. A famous example is the ouzo-effect, occurring in alcohols such as ouzo, pastis, and limoncello [19, 20, 21].
In biochemistry, phase separation has recently received a new wave of enthusiasm caused by the discovery of membrane-less compartments in the cell cytoplasm [22]. These compartments have been successfully described as droplets formed via phase separation [23] and are associated with various essential cell functions [24, 25]. Furthermore, alterations of membrane-less compartments’ physiological properties can trigger the formation of pathological protein aggregates [26]. Meanwhile, the fascinating idea that phase separation could have played a role in the primordial soup, originally proposed by Oparin and Haldane [27, 28], got reignited, thereby leading to new studies on mixtures composed of short oligomers [29, 30].
Building upon the recent theoretical and experimental advances in the thermodynamics of mixtures, in this paper we investigate the phase behavior of Cacio e pepe pasta sauce. Cacio e pepe (literally “cheese and pepper”) is a traditional recipe from Lazio, a region in central Italy. It consists of tonnarelli noodles served in a cream of pecorino cheese, pepper, and starch-enriched water, see Fig. 1(a). A precursor to this dish can be traced back to the fifteenth century, as documented in [31]. The modern version, however, is commonly linked to the long journeys of shepherds, who had to fill their saddlebags with highly caloric and easily transportable ingredients to sustain them during their travels. Pecorino cheese was ideal due to its extraordinary shelf life, black pepper was used to stimulate heat receptors, and homemade noodles provided the carbohydrate intake.
@@
Despite the short list of ingredients, preparing this dish requires extra care. This is because cheese and water mixtures are notoriously prone to separation when heated, as seen for example in Swiss cheese fondue [32]. In the context of Cacio e pepe, this challenge has already been explored by Dario Bressanini [33]. The most delicate step is the mixing of starch-enriched water with grated cheese. One starts cooking pasta noodles in boiling water, letting them release starch, then extracts part of the water and starch solution. An essential procedure is to wait some time before mixing water and cheese, to let the water cool down. This is because, at high temperatures, cheese proteins can either form clumps upon denaturation or simply aggregate [34, 35], therefore ruining the sauce, see Fig. 1(b). But temperature is not the only physical parameter that must be carefully controlled: protein aggregation and denaturation are concentration-dependent processes, thus, mixing the right amount of cheese, water, and starch is essential to avoid protein aggregates. In the absence of starch, for example, cheese in hot water forms huge clumps at a temperature around $65\degree\text{C}$, leading to what we name the “Mozzarella Phase”, see the first row of Fig. 1(b). If cheese is mixed with pasta water, which contains small amounts of starch, the emergence of clumps is reduced, and large protein aggregates are found at higher temperatures, see Fig. 1(b), second row. Pasta water can be “risottata”, i.e. collected and heated in a pan, so that some water evaporates and the starch is concentrated. If cheese is mixed with pasta water “risottata”, the presence of clumps is almost negligible, see the last row in Fig. 1(b).
@@
To overcome clump formation and achieve the perfect Cacio e pepe, in this work, we characterize the phase behavior of the solution containing water, starch, and cheese. We achieve this using common kitchen tools, ensuring that our results are easily reproducible not only by scientists in the lab but also by culinary enthusiasts. The paper is organized as follows: in Sec. II, we introduce the experimental setup and discuss the role of starch in the mixture phase behavior. In Sec. III, we fix the starch percentage and vary the cheese amount. Sec. IV discusses a minimal theoretical model that recapitulates our experimental finding. In Sec. V, we also investigate the role of a less traditional stabilizer, trisodium citrate, in preventing clump formation, analyzing its impact on sauce stability and comparing its effects to those of starch. Finally, in Sec. VI, we propose a recipe for the perfect Cacio e pepe.
II Increasing starch mitigates the formation of protein aggregates
@@
Figure 2: Starch mitigates protein clumps. (a) Phase diagram of the sauce state as a function of the starch percentage with respect to the water and temperature in Celsius. Each box contains a snapshot of the sauce mixture taken during the experiment, and its contour reflects the mean size of the corresponding sample via the color map shown on the left. The Mozzarella Phase indicates a region of the phase diagram where the cheese in the sample forms a clump of a size comparable to the one of the system. (b) The same phase diagram from (a) after applying Gaussian smoothing to better visualize phase behavior. The color represents the mean aggregate size in the sample. (c) Kernel regression smoothing of the phase diagram in a) to obtain a continuous map. The color map on the right refers to panels b) and c).
@@
Figure 3: Protein amount tunes aggregate formation. (a) Phase diagram of the sauce made by combining $50$ g of cheese with a varying amount of starch-enriched water (here expressed as a mass percentage) as a function of the mixture temperature. The starch percentage is fixed to $1$. Each box is colored with respect to the mean size of cheese clumps. (b) Same phase diagram as in (a) after Gaussian smoothing, where each point is colored according to the map of the mean aggregate size (color map on the right). (c) Phase diagram expressed as a function of protein mass fraction $\phi$, which is the component that leads to aggregation. Kernel regression smoothing has been applied to obtain a continuous diagram. The dashed gray line indicates the isoline of median aggregate size, while the dashed black line represents the parabolic fit separating homogeneous from clumped domains.
As discussed above, one of the most problematic aspects of making Cacio e pepe is the appearance of large aggregates of cheese (Mozzarella Phase). The presence of starch is the main difference between pasta water and normal water, suggesting that starch concentration is an important factor in determining the formation of cheese clumps.
In order to quantitatively investigate the role of starch, we conducted experiments at a fixed cheese-to-water ratio while varying the starch concentration in the water and the temperature of the system. To ensure quality and reproducibility and to comply with the traditional recipe, we used a Pecorino Romano DOP, whose nutritional values and certification details are discussed in the supplementary material, Sec. S1 and Table S1.
We prepared the starch-enriched water, dissolving dry corn starch into water at ambient temperature, targeting the weight percentages shown on the bottom axis of Fig. 2(a). The mixture was heated on a stovetop to gelatinization, a transition marked by a noticeable increase in viscosity and opacity. Afterward, we allowed the starch-enriched water to cool to room temperature to prevent excessive heating of the cheese-water mixture during blending. Cheese and starch-enriched water were combined in equal weights and thoroughly blended with a mixer. The resulting mixture was placed in a controlled heat bath, namely a pot of water maintained at a constant temperature using a sous vide cooker device. The mixture’s temperature was further monitored with an external thermometer positioned inside it, and samples were collected upon reaching thermal equilibrium, as indicated by a stable mixture temperature. Each sample was spooned onto a Petri dish and photographed. In the supplementary material, Sec. S1, we report on the experimental apparatus built to heat the mixture and collect images, along with a detailed explanation of the measurement protocol. Each mixture was imaged systematically increasing the heat bath temperature by $5$C at each sampling and the evaporated water was replenished at each temperature step. Experiments were performed at different starch percentages to generate the images forming the phase diagram in Fig. 2(a). At low temperatures, we did not observe protein aggregates independently of the starch concentration. However, by increasing the temperature, clumps start being detected as dark areas in the pictures in Fig. 2(a), thanks to a light source below the Petri dish (see Figs. S1 and S2 in supplementary material). To characterize the phase behavior of the mixture and quantify the presence of aggregates, we detected cheese clumps using quantile thresholding and approximated them as ellipses (see Sec. S2 in the supplementary material for details on data analysis). Since cheese aggregation results in a stringy sauce, a phenomenon often referred to in Italian culinary jargon as “filatura”, we quantified aggregate sizes by measuring the major axis of the corresponding ellipses, which we refer to as aggregate size. We note that other measures for characterizing aggregate (or relative) size yielded qualitatively similar results, as shown in Fig. S6 in the supplementary material. The mean aggregate size serves as the order parameter.
Samples with larger, well-separated clumps resulted in higher mean aggregate sizes, while smoother mixtures corresponded to lower values. From these measurements, we constructed a phase diagram by plotting the mean aggregate size as a function of starch concentration and temperature (Fig. 2(a)). We used a colored frame to visually associate each image with its corresponding order parameter value, measured in millimeters. To enhance interpretability, the phase diagram was smoothed using a Gaussian filter at discrete data points (Fig. 2(b)) and further refined with kernel regression smoothing to obtain a continuous representation (Fig. 2(c)). We excluded the samples in the Mozzarella Phase from the colormap, as their aggregate sizes are comparable to the system size and would distort the visualization. The final results reveal a clear quantitative impact of starch concentration: higher starch contents shift the transition to clumpier, less homogeneous mixtures to higher temperatures. Moreover, mixtures with low starch concentrations exhibited larger aggregate sizes, while higher starch contents reduced aggregate size, and eventually prevented the occurrence of the Mozzarella Phase.
Overall, these experiments corroborate the culinary insight that starch in pasta water stabilizes Cacio e pepe sauce providing a quantitative picture of its role in the cheese-water mixture’s phase behavior. Starch mitigates aggregation, shifts the onset of clump formation to higher temperatures, and decreases the size of aggregates, making the sauce less sensitive to temperature control errors during preparation.
III Cheese concentration tunes the emergence of protein aggregates
@@
Having described the sauce properties as a function of the starch, we ought to find the mixture’s degree of separation by varying the respective percentages of water and cheese. To this end, we fixed the percentage of starch in the water to $1$, a value potentially enabling the appearance of rich phase behavior. Following the procedure outlined above, we performed experiments at different percentages of starch-enriched water with respect to cheese. In Fig. 3(a), we show the resulting images, implementing a colored frame to highlight the mean size, namely the parameter employed to quantify the degree of separation. As expected, by fixing the concentration of starch to 1 %, we avoid the Mozzarella Phase emerging in some of the previous experiments (see Fig. 2). The smoothed phase diagram in Fig. 3(b) hints at the presence of a binodal region separating the homogeneous mixture from a domain where larger aggregates appear. Since proteins drive clump formation, it is natural to express the phase diagram as a function of their mass fraction, $\phi$, which we estimate from the cheese concentration using the nutritional values of Pecorino Romano DOP (see Table S1 in the supplementary material). Applying a kernel regression smoothing finally leads to a continuous phase diagram as a function of $\phi$ and $T$ (see Fig. 3(c)), which will also prove useful in the following modeling section. We then define the binodal line separating the aggregate-rich domain from the homogeneous one as the isoline of mean aggregate size, taken as the midpoint between the maximum and minimum observed values after smoothing. The shape of this binodal is a paraboloid that can be fitted with a simple quadratic functional form. The minimum of the parabola lies slightly below the protein mass fraction value $0.134$, achieved with water and cheese in 1:1 proportion. A parabola with positive curvature signals that both lower and higher values of protein mass fraction correspond to well-mixed sauces even at high temperatures. Finally, we report that in the phase space lying above the binodal (see Figs. 2 and 3) the clumps are dynamically stable and seem not to coarsen. This observation might suggest a Turing-like mechanism, where clump formation is governed by diffusion and active chemical reactions (i.e., reactions maintained out of equilibrium by continuous replenishment of chemical fuel) [36]. However, no such reactions take place in our emulsions. Instead, we believe that the lack of coarsening during experiments is simply due to the extremely slow ripening dynamics [37]
Figure 4: A minimal model for the phase behavior of Cacio e pepe sauce. (a) Cheese is composed of casein organized in micelles together with calcium phosphate and a small percentage of whey proteins. Upon heating, whey proteins denaturate, reaching a state that favors whey-whey and whey-casein aggregation. Furthermore, casein micelles aggregate on heating. (b) Interaction and relative size (respectively $\chi$ and $n$ in the effective free energy) obtained from experiments by solving Eq. (1) for different temperatures. (c) Comparison between theoretical and experimental binodal curve.